# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v1

@@ 3 Learning with arbitrary effective costs
In this mechanism, each agent $i$ starts with some initial choices of $s_{i}^{0}$ and shares that with the center. [^1] Note that, only the fraction $s_{i}$ is shared with the center and not the data, which is consistent with the principle of federated learning. The center also starts with a $w^{0}$ and broadcasts that and the entire initial data-contribution vector $s^{0}$ with all the agents. In every subsequent iteration $t$, each agent $i$ locally computes two quantities: (i) her updated contribution $s_{i}^{t+1}$ by taking one gradient ascent step w.r.t. her own contribution $s_{i}$, and (ii) the local gradient of agent $i$’s accuracy component of the social welfare w.r.t. $w$. Both are evaluated at the current values of $w^{t}$ and $s^{t}$ and sent back to the center. The center calculates an updated $w^{t+1}$ that averages (in spirit of the FedAvg algorithm [23]) all the local gradients sent by the agents. The center then shares $w^{t+1}$ and $s^{t+1}$ with all the agents. Formally, the updates are given as follows.
@@
The mechanism is detailed out in Algorithm 1. Our main result of this section is that under certain concavity and bounded derivative conditions, Algorithm 1 always converges to a Nash equilibrium. We need a few matrices, defined as follows for every $w$ and $s$. For $i,j\in N$ and $k,\ell\in\{1,\ldots,m\}$,
@@ 7             $s^{t+1}=s^{t}+\gamma g(w^{t},s^{t},\mu^{t})$
8             Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>a_{i}(w^{t},s_{i}^{t+1},s_{-i}^{t})$
@@ 9             Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
Remark. Note that the choice of step sizes $\gamma$ and $\nu$ scale inversely with $m$ and $n$ respectively. Moreover, we impose certain restrictions on $\lambda$ and $\tilde{\lambda}$, which may be restrictive. In Section 4, we remove such issue and propose a learning algorithm with choices of $\gamma$ and $\nu$ that do not scale with $m$ or $n$.
@@ 14             Send $d_{i}^{t+1}$ to the center
In the second phase of the algorithm, we update the model parameter $w^{t}$. Note that since all the agents contribute the maximum amount of data they own, this phase is simply pure federated learning (without incentive design). As such, each agent now computes the local gradient $d_{i}^{t+1}$ and sends it to the center. The center then aggregates the gradients and take a gradient ascent step with learning rate $\eta$. The center then broadcasts the updated parameter to the agents and the process continues.
@@
Suppose Assumption 3 holds and we consider the utility function with payment scheme given in Equation 8 with $\beta>\max_{i\in N}c_{i}^{\prime}(s_{i}^{\max})$. Then, for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$, 2P-UPBReD (Algorithm 2) converges to the Nash equilibrium $(w^{\texttt{{OPT}}},s^{\max})$ and is budget balanced. In particular, for any given $\epsilon>0$, we have
@@
where, $f(w,s):=-\frac{1}{n}\sum_{i\in N}a_{i}(w,s)$ and $\Delta_{i}=\beta-c_{i}^{\prime}(s_{i}^{\max})$.
@@ B.2 Proof of Theorem 2
Phase 1. Note that the derivatives of the utility of agent $i\in N$ is given by
$\begin{split}[g(w^{0},s^{t},\mu^{t})]_{i}&=\frac{\partial}{\partial s_{i}}u_{i}(w^{0},s^{t})\\ &=\frac{\partial}{\partial s_{i}}a_{i}(w^{0},s^{t})-\frac{\partial}{\partial s_{i}}c_{i}(s_{i})+\beta\\ &\geqslant\frac{\partial}{\partial s_{i}}a_{i}(w^{0},s^{t})-\frac{\partial}{\partial s_{i}}c_{i}(s_{i}^{\max})+\beta.\end{split}$ (16)
The inequality follows since $c_{i}$ is convex and differentiable everywhere in $S_{i}$ and its derivative is a non-decreasing function of $s_{i}$. This implies that the maximum value of $c_{i}(s_{i})$ is reached at $s_{i}=s^{\max}$. For the choice of $\beta>\max_{i\in N}c_{i}^{\prime}(s_{i}^{\max})$, we get $\Delta_{i}:=\beta-c_{i}^{\prime}(s_{i}^{\max})>0$. Since, $a_{i}$ is non-decreasing in $s_{i}$, we get that the RHS of Equation 16 is positive. Hence
$[g(w^{0},s^{t},\mu^{t})]_{i}>0,$
implying that at every step of this phase, the utility function increases for every agent. Applying the update rule for $s$ given by Algorithm 2 in phase 1, we get
$s^{t+1}$ $=s^{t}+\gamma g(w^{0},s^{t},\mu^{t})\geqslant s^{t}+\frac{\partial}{\partial s_{i}}a_{i}(w^{t},s^{t})+\Delta_{i},$
@@
$s^{t}_{i}\geqslant s^{0}_{i}+\gamma(t\Delta_{i}+\sum_{k=0}^{t}\frac{\partial}{\partial s_{i}}a_{i}(w^{0},s^{k})).$
@@
$s^{t}_{i}$ $\geqslant s^{\max}_{i}-l_{i}+\gamma(t\Delta_{i}+\sum_{k=1}^{t}\frac{\partial}{\partial s_{i}}a_{i}(s^{k},w^{k}))$
$\geqslant s^{max}_{i}-l_{i}+\gamma t\Delta_{i}$
where non-decreasingness of $a_{i}$ w.r.t. $s_{i}$ is used. Substituting $t=\kappa$ where $\kappa\geqslant\frac{\displaystyle l_{i}}{\displaystyle\Delta_{i}\gamma}$, we obtain
$\begin{split}s^{\kappa}_{i}&\geqslant s^{\max}_{i}-k_{i}+\gamma\Delta_{i}(\frac{\displaystyle k_{i}}{\displaystyle\Delta_{i}\gamma})\\ &\geqslant s^{\max}_{i}-k_{i}+k_{i}\\ &\geqslant s^{\max}_{i}\\ s_{i}^{\kappa}&\geqslant s_{i}^{\max}\end{split}$ (17)

### v2

@@ 3 Learning with arbitrary effective costs
In this mechanism, each agent $i$ starts with some initial choices of $s_{i}^{0}$ and shares that with the center. [^1] Note that, only the number of data points, $s_{i}$ is shared with the center and not the data, which is consistent with the principle of federated learning. The center also starts with a $w^{0}$ and broadcasts that and the entire initial data-contribution vector $s^{0}$ with all the agents. In every subsequent iteration $t$, each agent $i$ locally computes two quantities: (i) her updated contribution $s_{i}^{t+1}$ by taking one gradient ascent step w.r.t. her own contribution $s_{i}$, and (ii) the local gradient of agent $i$’s accuracy component of the social welfare w.r.t. $w$. Both are evaluated at the current values of $w^{t}$ and $s^{t}$ and sent back to the center. The center calculates an updated $w^{t+1}$ that averages (in spirit of the FedAvg algorithm [23]) all the local gradients sent by the agents. The center then shares $w^{t+1}$ and $s^{t+1}$ with all the agents. Formally, the updates are given as follows.
@@
The mechanism is detailed out in Algorithm 1. Our main result of this section is that under certain bounded derivative conditions, Algorithm 1 always converges to a Nash equilibrium. We need a few matrices, defined as follows for every $w$ and $s$. For $i,j\in N$ and $k,\ell\in\{1,\ldots,m\}$,
@@ 7             $s^{t+1}=s^{t}+\gamma g(w^{t},s^{t},\mu^{t})$
8             Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>a_{i}(w^{t},s_{i}^{t},s_{-i}^{t})$
@@ 9             Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
Remark. Note that the choice of step sizes $\gamma$ and $\nu$ scale inversely with $m$ and $n$ respectively. Moreover, we impose certain restrictions on $\lambda$ and $\tilde{\lambda}$, which may be restrictive. In Section 4, we remove such issues and propose a learning algorithm with choices of $\gamma$ and $\nu$ that do not scale with $m$ or $n$.
@@ 14             Send $d_{i}^{t+1}$ to the center
In the second phase of the algorithm, we update the model parameter $w^{t}$. Note that since all the agents contribute the maximum amount of data they own, this phase is simply pure federated learning (without incentive design). As such, each agent now computes the local gradient $d_{i}^{t+1}$ and sends it to the center. The center then aggregates the gradients and takes a gradient ascent step with learning rate $\eta$. The center then broadcasts the updated parameter to the agents and the process continues.
@@
Suppose Assumptions  2 and 3 hold and we consider the utility function with payment scheme given in Equation 8 with $\beta>\zeta+\tau$. Then, for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$, 2P-UPBReD (Algorithm 2) converges to the Nash equilibrium $(w^{\texttt{{OPT}}},s^{\max})$ and is budget balanced. In particular, for any given $\epsilon>0$, we have
@@
where, $f(w,s):=-\frac{1}{n}\sum_{i\in N}a_{i}(w,s)$ and $\Delta_{i}=\beta-\tau-\zeta$.
@@ B.2 Proof of Theorem 2
Phase 1. Using 1
$\begin{split}[g(w^{0},s^{t},\mu^{t})]_{i}\geqslant-\tau-\zeta+\beta>0\end{split}$ (16)
For the choice of $\beta>\zeta+\tau$ , we get $\Delta_{i}:=\beta-\zeta-\tau>0$. Applying the update rule for $s$ given by Algorithm 2 in phase 1, we get
$s^{t+1}$ $=s^{t}+\gamma g(w^{0},s^{t},\mu^{t})\geqslant s^{t}+\gamma\Delta_{i},$
@@
$s^{t}_{i}\geqslant s^{0}_{i}+\gamma(t\Delta_{i}).$
@@
$s^{t}_{i}$ $\geqslant s^{\max}_{i}-l_{i}+\gamma(t\Delta_{i})$
Substituting $t=\kappa$ where $\kappa\geqslant\frac{\displaystyle l_{i}}{\displaystyle\Delta_{i}\gamma}$, we obtain
$\begin{split}s^{\kappa}_{i}&\geqslant s^{\max}_{i}-l_{i}+\gamma\Delta_{i}(\frac{\displaystyle l_{i}}{\displaystyle\Delta_{i}\gamma})\\ &\geqslant s^{\max}_{i}-l_{i}+l_{i}\\ &\geqslant s^{\max}_{i}\\ s_{i}^{\kappa}&\geqslant s_{i}^{\max}\end{split}$ (17)